With a magnetic flux of 30,000 mx through a area o
Answer #1

62)
number of water molecules = 1.81 * 1024
number of atoms in water a molecule = 2*1+1 =3
total number of atoms in the sample = 3 * 1.81 * 1024 = 5.43 * 1024
number of moles of atoms in the given sample = 5.43 * 1024/NA = (5.43 *1024)/(6.022*1023) = 9.02